how much simple interest will be earned in 18 months form $750 placed in a savings account at 7.5% simple interest?
Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


The formula for simple interest is I = P * r * t where:


I = interest

r = yearly interest rate. If not time frame is given for the interest rate, such as "annual" or "daily" or "monthly", then yearly is assumed.

t = t in years. 18 months is 1.5 years.


In this case the I, interest earned is:

I = 750*(7.5/100)*1.5 = 84.375 rounded to $84.38